Contribution Margin and Contribution Margin Ratio

For a recent year, Wicker Company-owned restaurants had the following sales and expenses (in millions):

Sales $29,500
Food and packaging $10,055
Payroll 7,400
Occupancy (rent, depreciation, etc.) 6,865
General, selling, and administrative expenses 4,300
$28,620
Income from operations $880

Assume that the variable costs consist of food and packaging, payroll, and 40% of the general, selling, and administrative expenses.

a. What is Wicker Company's contribution margin? Round to the nearest million. (Give answer in millions of dollars.)
$ million

b. What is Wicker Company's contribution margin ratio? Round to one decimal place.
%

c. How much would income from operations increase if same-store sales increased by $1,800 million for the coming year, with no change in the contribution margin ratio or fixed costs? Round your answer to the closest million.
$ million
Answer #1

Variable cost = 10055+7400+(4300*40%) = 19175

a) contribution margin = Sales-Variable cost = 29500-19175 = 10325

b) Contribution margin ratio = 10325/29500 = 35%

c) Income from operation increase by = (1800*35%) = 630 million